OpenStudy (anonymous):

-q-11= 2q + 4 find q .

OpenStudy (anonymous):

This is simple linear eaquation -q-11=2q+4 -11+4=2q+q -7=4q -7/4=q just take the unknown term on one side and all the known terms on one side and solve that is the only way you can solve an linear equation

OpenStudy (anonymous):